OpenStudy (anonymous):

What is the solution of the equation? square root of x plus ten–9 = –7

OpenStudy (anonymous):

Could you re-write the equation with digits? For square root of x do root(x).

OpenStudy (whpalmer4):

\[\sqrt{x+10} - 9 = -7\]Add 9 to both sides \[\sqrt{x+10} = 2\]Square both sides and solve for x.

OpenStudy (anonymous):

The answer choices are –6 4 14 –8

OpenStudy (anonymous):

Is WhPalmer's equation what you're working with?

OpenStudy (anonymous):

That is, what is under the square root symbol?

OpenStudy (whpalmer4):

The answer to the problem I wrote is in the list of choices...

OpenStudy (whpalmer4):

And I think it's the only problem that makes sense. There has to be something on the left side large enough to make -7 positive if the item from the left moves to the right; otherwise, we're talking square roots of a negative number, and I doubt that is the case.

OpenStudy (whpalmer4):

@carlycarl3 do you know what would be on the left hand side if you squared it?

OpenStudy (whpalmer4):

Here's a hint: \[(\sqrt{z})^2 = z\]